OpenStudy (anonymous):

What is the exact area of a circle with a radius of 12 feet? A. 24π square feet B. 36π square feet C. 121π square feet D. 144π square feet

OpenStudy (johnweldon1993):

The area of a circle is found by the formula \[\large A = \pi r^2\] So you can plug in your 'r' (radius) and you will have your exact Area

OpenStudy (anonymous):

\[A = \pi r ^{2}\]

OpenStudy (anonymous):

Well it seems i was a little slow on the trigger

OpenStudy (anonymous):

im lost i dont know but i got 75.36

OpenStudy (anonymous):

the answer is 144

OpenStudy (anonymous):

Read the answers you dont have to multiply by Pi to know which is right

OpenStudy (anonymous):

And no its not 144 read the answers you can select

OpenStudy (johnweldon1993):

That is because you evaluated...thats exactly why they put EXACT area...we know \(\large \pi \) to be an irrational number....so it can never be EXACT...just estimated... So when we have \[\large A = \pi r^2\] and the R is 12..we have \[\large A = \pi (12)^2\] and we just do the 12^2 to be 144 so we have \[\large A = 144\pi\]

OpenStudy (anonymous):

the formula of area of a circle is pir^2 - so than will be area =144pi

OpenStudy (anonymous):

Oh i apologize i misthought i read it as r2 not r^2

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

What is the exact area of a circle with a diameter of ft? A.1936 pi square feet B.484 pi square feet C.88 pi square feet D.44 pi square feet

OpenStudy (anonymous):

i do believe it is B

OpenStudy (anonymous):

What is the approximate area of a circle with a radius of 28 centimeters? Use π ≈ 3.14. A. 2461.8 square centimeters B. 87.9 square centimeters C. 175.8 square centimeters D. 4923.5 square centimeters